OpenStudy (anonymous):

Solve for b. a = a+b+c/3 A.) 3A - a - c = b B.) -c / 3 = b C.) A - c / 3 = b C.) -3 - c = b

OpenStudy (anonymous):

letter A.

OpenStudy (ash2326):

We have \[a=a+b+\frac c3\] Subtract both sides by a \[a-a=a-a+b+\frac c3\] We get \[0=0+b+\frac c3\] now subtract both sides by c/3 \[0-\frac c3=b+\frac c3-\frac c3\] we get finally \[-\frac c3=b\]

OpenStudy (anonymous):

a=a+b+c/3 3a-a-c=b

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

Pass the a that is in the denominator multiplying the letter "a"